1. Show that the atomic packing factor for FCC is 0.74.
2. Show that the atomic packing factor for BCC is 0.68.
3. Iron has a BCC crystal structure, an atomic radius of 0.124 nm, and an atomic weight of 55.85 g/mol. Compute and compare its theoretical density with the experimental value, 7.87 g/cm3.
4. Rhodium has an atomic radius of 0.1345 nm and density of 12.41 g/cm3. Determine whether it has an FCC or BCC crystal structure.
5. The unit cell for tin has tetragonal symmetry, with a and b lattice parameters of 0.583 and 0.318 nm, respectively. If the density, atomic weight, and atomic radius are 7.30 g/cm3, 118.69 g/mol, and 0.151 nm, respectively, compute the atomic packing factor.
1
Expert's answer
2020-12-01T09:28:39-0500

1. "APF" is the fraction of solid sphere volume in unit cell.

"APF=" "Total volume of sphere\over Total unit cell volume" "=" "V_s\over V_c"

"V_s=(4)." "4\over3" "\Pi R^3=" "16\over 3" "\Pi R^3"

"V_c=" "16R^3" "\sqrt{2}"

"APF=" "V_s\over V_c" "=" "{{16\over3}\Pi R^3\over 16R^3\sqrt{2}}" "=0.74"

2. "BCC=0.68"

In terms of "a\infin" "R"

Atoms per unit cell

"a=" "4\over \sqrt{3R}"

Coordination "=no.8"

"APF=0.68"

3. In finding the density of "Fe"

"\rho=" "_nA_{Fe}\over V_cN_A"

For "BCC, n=2" atoms /unit cell, and

"V_c=" "(" "4R\over V3" ")^3"

"\rho=" "_nA_{Fe}\over {({4R\over V3})}^3N_A"

"{(2atoms/unit)(55.85g/mol)}\over[(4)(0.124\times 10^{-7}/unit cell(6.023\times10^{23}atoms/mol)]" "=7.90g\/cm^3"

The actual value of "Fe=7.87g\/cm^3"

4. "Rd=0.1345nm"

"\rho=12.41g/cm^3"

"A_w=102.91g\/mol"

"\rho=" "nA_w\over V_cN_A"

"V_{c_{BCC}}=a^3_{BCC}=" "{({4R\over V3})}^3" "n_{BCC}=2"

"V_{c_{BCC}}=a^3_{FCC}=(2R\sqrt{2})^3" "n_{FCC}=4"

For FCC, "\rho=" "4(102.91g/mol)\over{{({(4)\times (0.1345\times 10^{-7}cm^3\over \sqrt{3}})}}" "\times 6.023\times 10^23atoms/mol"

"=11.41g\/cm^3"

It is "FCC"

5. "a=0.583nm"

"b=0.318nm"

"\rho=7.30g/cm^3"

"A_w=118.69g\/mol"

"A_R=0.151nm"

"n=" "\rho V_cN_A\over A_{sn}"

"=(7.3g/cm^3)(5.83^2)(3.18\times 10^{-24})(6.023\times 10^{23}atoms/mol)"

"=4atoms\/ unit cell"

"\therefore" "APF=" "V_s\over V_c" "=" "(4)({4\over 3}\Pi R^3)\over(a)^2(b)"

"=" "4({4\over 3}\Pi(1.52\times 10^{-8}cm^3))\over(a5.83\times 10^{-8}cm^2)(3.18\times 10^{-8})"

"=0.544"